Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.
Find the total cost of leaves as percent of gross salary.
1
Expert's answer
2019-05-22T08:55:59-0400

Gross salary = 18000 + 9000 = 27,000 Rs. 

Cost of casual leaves per year = {12 / (22 x 12)} x 27000 x 12 = 14,727.27 Rs. 

Cost of earned leaves per year = {24 / (22 x 12)} x 27000 x 12= 29,454.55 Rs. 

Total cost of leaves per year = 14,727.27 + 29,454.55 = 44,181.82Rs. 


Total cost of leaves as percent of gross salary


"{44181.82 \over 12\times 27000}\times100\%=13.64\%"